Two bonds are given.

  1. A 9.3%, 23-year semiannual bond at 112.429
  2. A 11.4%, 26-year semiannual bond at 100.834

Assume that the par value of the bond is $1,000.

  1. Which of these two bonds offers the highest current yield?
    -Select- Item 1 or 2
  2. Which one has the highest yield to maturity?
    -Select Item 1 or 2
Answer #1

1.
Bond 2 offers the highest current yield
Current Yield=Coupon rate*Face Value/Price
Bond 1=9.3%*100/112.429
=8.2719%

Bond 2=11.4%*100/100.834
=11.3057%

2.
Bond 2 offers the highest yield to maturity
Bond 1=RATE(23*2,9.3%*100/2,-112.429,100)*2=8.1000%

Bond 2=RATE(26*2,11.4%*100/2,-100.834,100)*2=11.3000%